An insulated rigid tank is divided into two equal parts by a partition. Initially, one part contains 15 kg of water vapor with steam quality 20% at 20 bar and the other part is evacuated. The partition is removed, and the water vapor expands into the entire tank.a) Determine the final temperature and pressure using higher accuracy calculation. b) Recalculate (a) by assuming the gas behave as ideal gas
Answer #1

Any LLL 15 Hy evacuated 2=0. p=20 bar luated From saturated water pressure table at 20 bar (7-5) T59+ = Ti = 212.38C Ve = 0.. at UF Uft a vez = 5045+012 (2024.6) gog.42 e V=0.356 m3kg water table. From Saturated T = 102°C Pa = 111 - PQI as gdeal ga